### v1

# Domain Specific Pruning of Large Mixture-of-Experts Models with Few-shot Demonstrations

Abstract
Mixture-of-Experts (MoE) models achieve a favorable trade-off between performance and inference efficiency by activating only a subset of experts. However, the memory overhead of storing all experts remains a major limitation, especially in large-scale MoE models such as DeepSeek-R1 (671B). In this study, we investigate domain specialization and expert redundancy in large-scale MoE models and uncover a consistent behavior we term few-shot expert localization, with only a few demonstrations, the model consistently activates a sparse and stable subset of experts. Building on this observation, we propose a simple yet effective pruning framework, EASY-EP, that leverages a few domain-specific demonstrations to identify and retain only the most relevant experts. EASY-EP comprises two key components: output-aware expert importance assessment and expert-level token contribution estimation. The former evaluates the importance of each expert for the current token by considering the gating scores and magnitudes of the outputs of activated experts, while the latter assesses the contribution of tokens based on representation similarities after and before routed experts. Experiments show that our method can achieve comparable performances and 2.99×2.$9$ throughput under the same memory budget with full DeepSeek-R1 with only half the experts. Our code is available at https://github.com/RUCAIBox/EASYEP.
Figure 1: Throughput and Performance Comparison of DeepSeek-R1 on AIME2024 with Varying Expert Counts Using EASY-Prune. We deploy R1 with two 8$\times$ H800 for 224 and 256 experts while one 8×8$\times$8 × H800 is employed for others. The throughputs of latter configurations are multiplied by 2.
1 Introduction
Mixture-of-Experts (MoE) architectures have been widely adopted as the backbones of various large language models (LLMs) due to their efficiency of scaling parameters without proportional computational overhead [1, 2, 3, 4]. However, the deployment of large MoE models imposes substantial memory requirements. Taking DeepSeek-R1 (671B) [1] as an example, it takes about 1500 GB under BF16 precision and 750 GB under FP8 precision, necessitating 4×8 A800 or 2×8 H800 GPU configurations, respectively. This underscores the critical need to explore lite deployment strategies for large-scale MoE models like DeepSeek-R1.
Various training-free approaches have been proposed to alleviate the inference memory demands of MoE models. Router-based pruning methods leverage statistical information from router activation and gating values to estimate expert importance [5], while perturbation-based strategies search for the optimal subsets of experts to minimize the distribution drift of hidden states [6, 7]. Despite these advancements, achieving an optimal trade-off between efficiency and performance remains a challenge. Router-based methods often fail to accurately identify critical experts, whereas perturbation-based approaches become computationally prohibitive as the number of experts scales up. Moreover, existing methods are primarily designed for MoEs with a few experts per layer (e.g., Mixtral 8$\times$7B [2][6]. This highlights the need for more scalable and accurate pruning strategies tailored to large-scale MoE models. While such a scale poses challenges for current approaches, it also brings new opportunities: the increased granularity and domain specialization of experts in models like DeepSeek-R1 [8, 9] make them especially amenable to domain-specific pruning at high compression ratios [10].
In this study, we analyze how expert activations in a large MoE model vary across domains and respond to a small number of demonstration samples. We observe a consistent behavior: given just a few domain-specific examples, the model tends to activate a stable and sparse subset of experts that are highly relevant to that domain. We refer to this phenomenon as few-shot expert identification. To better understand this behavior, we first examine domain-level differences in expert activation distributions and the role of high gating-value experts. We then analyze how the number of demonstrations and the similarity across datasets within the same domain affect expert identification. Our key findings are as follows: (1) High gating-value experts are strongly domain-specific, which play a critical role in their respective domains but have minimal or no influence on unrelated tasks; (2) A small number of demonstrations is sufficient to identify these experts reliably, and the identified experts generalize well to other datasets within the same domain.
Based on our observations, we introduce a domain-specific pruning framework that leverages few-shot demonstrations to address memory constraints in large-scale MoE models. Our approach begins by sampling a small number of task demonstrations from a specific domain and generating responses with the original model, serving as a calibration set. We then introduce a pruning method, Expert Assessment with Simple Yet-effective scoring for Expert Pruning, a.k.a., EASY-EP. It aims to quantify expert importance and retain a fixed-size subset of top-scoring experts. This proposed metric consists of two complementary components: output-aware expert importance assessment and expert-level token contribution estimation. The former component jointly considers gating values and L2 Norm of expert outputs to evaluate per-token expert importance, while the latter component uses similarity between input and residual-connected output of routed experts to measure the contribution of individual tokens to the overall expert score. Thus, a single forward pass is sufficient to identify target domain-specific experts without the need for backpropagation or repeated forward evaluations.
To evaluate the efficacy of our approach, we conducted systematic experiments on DeepSeek-R1 using six benchmark datasets spanning math, coding, science, and agent execution capabilities. Specifically, our method achieves $100.38\%$ and $95.10\%$ performance compared to the full model while retaining only $50\%$ of experts on domain-specific and mixed-domain pruning settings, respectively. As the compression ratio increases, the performances degrade slowly, indicating robustness to pruning. Additionally, under identical memory constraints, pruning $50\%$ of the experts leads to a 2.99$\times$ increase in inference throughput under the setting of 1K input and 1K output lengths, highlighting the practical utility of our framework for real-world deployments.
2 Background
Recently, Mixture-of-Experts (MoE) architectures have been widely adopted as the backbone of LLMs due to their ability to scale parameters efficiently without incurring additional computational costs [2, 10]. This architecture introduces MoE modules where parameters are dynamically activated to replace traditional feedforward networks (FFNs) In the $l$-th layer, a MoE module consists of a router $\operatorname{G}(\cdot)$, $N$ routed experts $\{\operatorname{E}^{l}_{1}(\cdot),\dots,\operatorname{E}^{l}_{N}(\cdot)\}$, and an optional shared expert $\operatorname{E}^{l}_{s}(\cdot)$  [^1] Shared experts were introduced by the DeepSeekMoE series [10, 1, 11, 12], but are not considered in our analysis and method. at layer $l$. Given an input representation sequence $\mathbf{H}^{l}=\{\bm{h}^{l}_{1},\dots,\bm{h}^{l}_{T}\},\forall\bm{h}_{t}^{l}\in\mathbb{R}^{D}$, the router computes the logits of each routed expert for the $t$-th token and applies a gating function on the Top-$K$ logits to obtain the gating values:
$g^{l}_{i,t}=\begin{cases}\operatorname{G}^{l}_{i}(\bm{h}_{t}^{l}),\quad&\operatorname{G}^{l}_{i}(\bm{h}_{t}^{l})\in\text{Top-}K({\operatorname{G}^{l}_{1}(\bm{h}_{t}^{l}),\dots,\operatorname{G}^{l}_{N}(\bm{h}_{t}^{l})}),\\ 0,&\text{otherwise}.\end{cases}$ (1)
Subsequently, the Top-$K$ selected experts are activated, and their outputs are aggregated via weighted summation. The final output is obtained by summing the input hidden states and the output of routed experts as follows:
$\tilde{\bm{h}}^{l}_{t}={\bm{h}}^{l}_{t}+\sum_{i=1}^{N}g^{l}_{i,t}\cdot\operatorname{E}_{i}^{l}(\bm{h}^{l}_{t})$ (2)
With the gating values of the router, we define two metrics to assess the importance of each expert, i.e., frequency and gating scores [5]. For all tokens in a calibration set and an expert $\operatorname{E}_{i}^{l}$, we define frequency $f^{l}_{i}$ as the number of times each expert is activated, while gating scores $r^{l}_{i}$ is defined as the total sum of the gating values when each expert is activated.
$f^{l}_{i}=\sum_{n=1}^{M}\sum_{t=1}^{T_{n}}(g^{l}_{i,n,t}>0),\quad r^{l}_{i}=\sum_{n=1}^{M}\sum_{t=1}^{T_{n}}g^{l}_{i,n,t},$ (3)
where $M$ is the number of demonstrations in the calibration set, and $T_{n}$ denotes the number of tokens per demonstration.
3 Empirical Analysis of Experts
Previous work has demonstrated that the expert distributions of MoEs with few experts (e.g., Mixtral $8\times 7$B) mainly depend on the syntax structures instead of tasks and domains of input data [2]. However, the large-size MoE models are equipped with various fine-grained experts, e.g., DeepSeek-R1 has 256 experts in each layer. Thus, the experts may be more specialized and store distinct knowledge and capacities in their parameters. To investigate this, we empirically investigate what we refer to as few-shot expert identification, a phenomenon in which domain-specific experts can be reliably identified using only a small number of demonstrations.
Figure 2: Overlap ratios of experts with top gating scores on different datasets.
3.1 Expert Specialization Across Domains
To investigate whether experts in large MoE models exhibit domain-specific specialization, we select AIME-2023  [^2] https://huggingface.co/datasets/di-zhang-fdu/AIME_1983_2024, GPQA-main [13], and LiveCodeBench-V3 [14] as representative datasets for the domains of math, science, and coding, respectively. We conduct experiments using the representative model DeepSeek-R1 on these datasets and extract the gating scores across different domains (as described in Section 2). By analyzing the expert activation distributions across these domains, we aim to reveal how expert utilization varies and assess the degree of specialization.
Distinct Expert Distribution Across Domains. We first rank all experts at each layer by their gating scores and select the top-$M$ experts for each dataset, where $M\in\{2,16,128\}$. We then measure the overlap of top-ranked experts across different domains and visualize how this overlap varies with $M$ in Figure 2. We observe that the top-2 experts are completely disjoint across datasets, and even as $M$ increases, a substantial number of non-overlapping experts remain. This indicates that large MoE models contain domain-specialized experts that are predominantly activated in their respective domains.
Table 1: Results of removing domain-specific experts.
Domain AIME24 GPQA LiveCodeBench
Full 78.00 70.91 63.32
Math 67.33 (-10.67) 69.19 (-1.72) 65.27 (+1.95)
Code 78.67 (+0.67) 71.72 (+0.81) 55.68 (-6.64)
Science 79.33 (+1.33) 59.09 (-11.82) 61.07 (-2.25)
Impact of Removing Domain-Specific Experts. In order to explore the importance of domain-specific experts, we remove those that appear among the top-128 by gating score in one domain but not in the top-128 of any other domain. We then evaluate each pruned model on tasks from the same domain as the calibration data (i.e., in-domain), as well as on tasks from other domains (i.e., out-of-domain), to assess the generalizability of the retained experts. As shown in Table 1, pruning these experts leads to significant performance degradation on in-domain tasks while having minimal impact on out-of-domain tasks. From these findings, we can conclude that domain-specific experts play a critical role in relevant domain but are redundant for other domains.
3.2 Expert Locality Within One Domain
Beyond examining the expert specialization across different domains, we further investigate the distribution of experts within the same domain. Specifically, we study how the number of demonstrations affects pruning and whether the identified critical experts remain highly activated on other datasets within the same domain.
Figure 3: Left: Overlap ratio of top-128 experts with different number of demonstrations. Right: Overlap ratio of top-128 experts pruned with different math datasets.
Effect of Calibration Set Size. Given the presence of domain-specific experts in DeepSeek-R1, a fundamental question arises: How many demonstrations are necessary to accurately identify these key experts? To investigate this, we sample varying numbers of demonstrations (i.e., 1,3,5,10,25,100) from LiveCodeBench-v3 [14]. For each setting, we compute the average gating scores and retain the top 128 experts. Based on these selections, we then calculate the pairwise overlap ratios between expert sets derived from different demonstration sizes. As illustrated in Figure 3 (Left), even with just five demonstrations, over $90\%$ of the critical experts can be effectively identified. Furthermore, 25 demonstrations are sufficient to capture all domain-specific experts ($99\%$), with additional demonstrations yielding only marginal improvements. These results underscore the feasibility of few-shot domain-specific expert pruning.
Consistency of Expert Activation Across Datasets. To investigate the consistency of domain-specific expert activation across datasets within the same domain, we conducted experiments with DeepSeek-R1 on four math datasets: AIME-2023, AIME-2024, AIME-2025, and HMMT-Feb 2025. Specifically, we perform expert pruning under the 25-shot setting for each dataset and retain the top 128 experts based on their average gating scores. We then compute the pairwise overlap ratios between each pair of datasets. As shown in Figure 3 (Right), the overlaps exceed 84% across all math datasets, revealing a high degree of consistency in domain-specific expert activation. This indicates that domain-specific expert activation patterns are largely transferable within the same domain. Despite some dataset-specific differences, the overall expert overlap remains strong and stable.
4 Method
4.1 Overview
Motivated by our earlier observation of few-shot expert identification phenomena in Section 3, we propose an expert pruning framework to reduce memory costs. Given a small set of demonstrations from the target domain, we run them through the MoE model and collect statistics on expert activation, considering both the inputs and outputs of each expert for pruning purposes. To effectively identify domain-specific experts in large MoE models, we introduce a simple yet effective expert pruning method, EASY-EP. Specifically, we first compute the output-aware expert importance $c_{i,t}^{l}$ as the product of the expert output L2 norm and its corresponding gating value.
Figure 4: Overall framework of EASY-EP. Given a calibration set consisting of input and responses by the model, EASY-EP leverages output-aware expert importance assessment and expert-level token contribution estimation to compute the expert score on the domain and returns the pruned expert sets.
Next, we determine the expert-level token contribution $s^{l}_{t}$ based on representation similarity between before and after expert computation. The final expert score $I(\operatorname{E}_{i}^{l})$ is obtained by multiplying these two terms:
$I(\operatorname{E}_{i}^{l})=\sum_{t=1}^{T}c_{i,t}^{l}\cdot s^{l}_{t}.$ (4)
Our method also supports mixed-domain pruning by averaging the normalized expert scores of all target domains. Thus, we can prune a single model to handle tasks across multiple domains:
$I_{mix}(\operatorname{E}_{i}^{l})=\sum_{\tau\in\mathcal{T}}(I_{\tau}(\operatorname{E}_{i}^{l})/\sum_{j=1}^{N}I_{\tau}(\operatorname{E}_{j}^{l})).$ (5)
Based on the expert scores computed from a small subset of data, we can efficiently select the Top-$M$ experts with the highest scores as retained experts while pruning other experts to reduce memory costs. The overall framework of our method is illustrated in Figure 4.
4.2 Output-Aware Expert Importance Assessment
To assess the importance of each expert, prior router-based expert pruning methods [5] assume activated expert gating scores can reflect their importance. However, this assumption has not been thoroughly examined. To analyze this, we begin by examining the aggregated output $\bar{\bm{h}}_{t}^{l}$ from all routed experts for a given token:
$\bar{\bm{h}}_{t}^{l}=\sum_{i=1}^{N}g_{i,t}^{l}\cdot\bm{e}_{i,t}^{l}=\sum_{i=1}^{N}g_{i,t}^{l}\bm{\lVert}\bm{e}_{i,t}^{l}\rVert\cdot\frac{\bm{e}_{i,t}^{l}}{\lVert\bm{e}_{i,t}^{l}\rVert},$ (6)
[…]
$\lVert\bar{\bm{h}}_{t}^{l}\rVert\leq\sum_{i=1}^{N}\Bigl{\lVert}g_{i,t}^{l}\bm{\lVert}\bm{e}_{i,t}^{l}\rVert\cdot\frac{\bm{e}_{i,t}^{l}}{\lVert\bm{e}_{i,t}^{l}\rVert}\Bigr{\rVert}=\sum_{i=1}^{N}g_{i,t}^{l}\bm{\lVert}\bm{e}_{i,t}^{l}\rVert.$ (7)
Figure 5: Left: Gating scores and averaged product of gating value and L2 Norm of expert outputs. Blue/red dots indicate experts in top-128 gating scores $g_{i,t}^{l}$; blue/yellow dots denote experts in top-128 expert importance $c_{i,t}^{l}$; black dots indicate neither. Right: Consine similarity between representations before and after incorporating the outputs of the routed expert. The red box and green box indicate high similarity and low similarity, respectively.
This indicates that the each expert’s contribution to the final output is bounded by the product of its gating value and the L2 norm of its output, $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert$. To empirically validate this insight, we visualize in Figure 5 (Left) the relationship between gating scores $g_{i,t}^{l}$ and the product $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert$. We can observe an expert with a large gating score may still produce outputs with low L2 Norm, ultimately resulting in a limited influence on the final output. Therefore, instead of using only the gating value, we define the importance of an expert for a given token as the product of its gating value and L2 Norm of output, as formalized in the following equation:
$c_{i,t}^{l}=g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert,\quad\forall g_{i,t}^{l}>0.$ (8)
4.3 Expert-Level Token Contribution Estimation
When calculating the statistical metrics for evaluating the importance of experts, prior work often directly averages the scores across all tokens [6, 5]. However, in practice, the influence of routed experts’ outputs on the residual stream varies across tokens. As illustrated in Figure 5 (Right), the similarity between representations before and after incorporating the outputs of routed experts differs significantly across tokens for DeepSeek-R1. For some tokens at specific layers, skipping the expert module results in minimal changes to the hidden states, with over 99% similarity between the input and output representations (e.g., the 15th layer of the first tokens). In contrast, for certain tokens and layers (e.g., the 12th layer of the first tokens), the hidden states show substantial differences after expert routing. Intuitively, for the latter type of tokens, modifying their routed experts could introduce a considerable distribution drift in representations [15]. This observation further suggests that such tokens should be given greater weight when assessing the importance of experts.
Inspired by these observations, we propose a similarity-based token importance assessment method. Given the representations before and after the routed expert modules $\bm{h}_{t}^{l}$ and $\tilde{\bm{h}}_{t}^{l}$, we compute the cosine similarity between these representations. The token importance score $s_{t}^{l}$ is then defined as one minus this similarity, capturing the extent of change induced by the routed expert module:
$s^{l}_{t}=1-\operatorname{Sim}(\bm{h}_{t}^{l},\tilde{\bm{h}}_{t}^{l}).$ (9)
5 Experiments
5.1 Experimental Settings
Evaluated Benchmark. To systematically assess the effectiveness of our proposed method, we conduct experiments across six benchmark datasets: AIME-2024, AIME-2025, HMMT-Feb 2025, LiveCodeBench [14], GPQA-Diamond [13], and AgentBench-OS [16]. These benchmarks encompass four fundamental domains and tasks of LLMs: math, coding, science, and agent-based task execution. We set the maximum context length to 32K, the temperature to 0.6, and the top-p sampling value to 0.95. To ensure statistical reliability, each benchmark is evaluated independently five times, and we report the average performances of pass@1.
Pruning Settings. To demonstrate the domain specialization among DeepSeek-R1’s experts, we perform domain-specific pruning and utilize distinct calibration sets for different domains. For each domain, we randomly sample 25 instances and construct a calibration set by concatenating their inputs with the corresponding outputs generated by DeepSeek-R1. We then evaluate the experts’ scores on the calibration data with our method and select the top 64 and 128 experts with the highest scores at each layer, respectively. We also average the normalized expert scores on different domains to evaluate the mixed-domain pruning performances. Details regarding the candidate sets for each domain are provided in Appendix A.
Baselines. In our experiments, we employ three baseline methods for comparison: (1) Random: A fixed number of experts are randomly selected, while the remaining experts are pruned. (2) Frequency: The Top-$M$ experts with the highest activation frequencies are retained. (3) Gating Score: The Top-$M$ experts with the highest gating scores are retained. The definitions of frequency and gating scores metrics are provided in Equation 3. We do not include perturbation-based methods in our comparison [7, 6], as they require evaluating changes in representations before and after pruning, which is computationally prohibitive for DeepSeek-R1 with 256 experts (the detailed analysis is shown in Appendix B).
5.2 Main Results
Table 2: Comparison of performances of different expert pruning methods. HMMT denotes HMMT-Feb 2025, GPQA denotes GPQA-Diamond, and A-OS denotes AgentBench-OS.
[…]
Full - 256 78.00 65.33 45.33 63.32 70.91 40.51 60.57 -
Random $\times$ 64 0.00 0.00 0.00 0.00 26.09 0.00 4.35 7.18
Frequency $\times$ 64 0.00 0.00 0.00 0.00 17.68 2.78 3.41 5.63
Gating Score $\times$ 64 2.67 1.33 2.67 0.00 20.20 0.69 4.59 7.58
EASY-EP $\times$ 64 72.67 55.33 36.00 42.51 67.47 27.26 50.20 82.88
Random $\times$ 128 8.33 6.67 3.33 20.96 41.41 7.64 14.72 24.30
Frequency $\times$ 128 19.33 13.33 7.33 16.17 60.01 29.16 24.22 39.99
Gating Score $\times$ 128 69.33 58.67 37.33 46.71 62.63 31.94 51.10 84.37
EASY-EP $\times$ 128 80.67 66.67 48.33 61.11 70.12 37.92 60.80 100.38
Frequency 128 18.00 13.33 8.00 9.58 30.30 13.19 15.40 25.42
Gating Score 128 18.67 8.67 8.67 53.29 37.37 25.00 25.28 41.73
EASY-EP 128 80.67 62.00 39.33 58.38 70.51 34.72 57.60 95.10
Table 2 shows the performance comparison of our method with baselines on different pruning settings. Firstly, our method outperforms other baselines and even surpasses the full DeepSeek-R1 model with only half of the experts on certain datasets for domain-specific pruning. Across all benchmarks and pruning settings, our approach consistently achieves significantly better results than all baselines. Moreover, under domain-specific pruning settings, our method maintains performance comparable to full DeepSeek-R1 and even exceeds it on math-related benchmarks. We hypothesize that removing irrelevant experts allows the model to rely more on domain-specific knowledge, leading to improved performance.
Secondly, our method can maintain performance even with a large compression ratio, as shown in Figure 1. When the compression ratio is large (e.g., $75\%$), most pruning methods can hardly correctly complete the query at all, manifested as 00 scores on many datasets. Additionally, they even fail to follow instructions correctly and produce nonsensical speech, e.g., failing to generate option letters for GPQA. On the contrary, our method can preserve $82.88\%$ averaged performances across all datasets and even $95.15\%$ performances on GPQA.
Finally, our method can also preserve performances under mixed-domain pruning. By leveraging mixed-domain pruning, our approach effectively preserves performance across multiple target domains, retaining $95.09\%$ of the original performance, which even surpasses domain-specific pruning methods via gating score metric. In contrast, other expert pruning techniques struggle to maintain a balanced performance across different domains. This result highlights that the overlapped experts identified by our method, which are associated with general reasoning abilities, can effectively contribute to a wide range of downstream domains.
5.3 Detailed Analysis
In this section, we conduct further ablation studies and detailed analyses to investigate the effectiveness of our approach and few-shot expert identification phenomena of DeepSeek-R1.
5.3.1 Ablation Study
We conduct ablation studies to analyze the impact of each component in our method. Specifically, we evaluate two variants with 64 remaining experts: (1) removing the token-level contribution estimation and (2) replacing the product of gating values and L2 Norm of expert outputs with only the gating scores. As shown in Table 3, both incorporating token-level contribution estimation and considering the L2 Norm of expert outputs lead to improved performance compared to using only the gating score. Furthermore, combining both components results in the best overall performance. These findings highlight the importance of both components in our method.
Table 3: Results of ablation study. Norm denotes whether considering L2 Norm of expert output and Token denotes whether considering token contribution scores.
Method Metric Experts AIME-24 AIME-25 HMMT LiveCode GPQA A-OS
Ours $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert\cdot s^{l}_{t}$ 64 72.67 55.33 36.00 42.51 67.47 27.26
w/o Token $g_{i,t}^{l}\lVert\bm{e}_{i,t}^{l}\rVert$ 64 65.33 49.33 31.33 27.54 56.57 21.53
w/o Norm $g_{i,t}^{l}\cdot s^{l}_{t}$ 64 70.00 40.00 23.33 19.76 61.11 18.75
w/o both $g_{i,t}^{l}$ 64 2.67 1.33 2.67 0.00 20.20 0.69
5.3.2 Effect of Pruning Data
We also explore the effect of pruning data. Instead of the input and output of full DeepSeek-R1, we consider full different data: (1) Inp: only input data; (2) R1: only generation data by DeepSeek-R1; (3) Inp+Ans: input data and golden answer; and (4) PT: pre-training data within the same domain. The experiment results are displayed in Table 4. Compared to using the concatenation of input and DeepSeek-R1’s output, the model’s performance under other settings suffered degradation in both math and coding tasks. This indicates that even within the same domain, distinct differences of expert distributions exist between the input data, golden answer, the model’s reasoning processes, and the pretraining data.
5.3.3 Generalization Capacity
In addition to evaluating performance on tasks within the same domain as the pruning data, we also assess tasks from unrelated domains (e.g., pruning DeepSeek-R1 with AIME2023 data while evaluating on LiveCodeBench). As presented in Table 5, out-of-domain performance declines compared to in-domain results but still retains a notable level of generalization capability. We hypothesize that although some experts associated with specific domain knowledge were pruned, experts related to core reasoning abilities were preserved. Specifically, the model demonstrates better generalization performance between math and science domains, which may be attributed to the inherent conceptual similarities between these fields.
Table 4: Comparison of performances with different pruning data.
Data AIME24 LiveCodebench GPQA
Inp+R1 80.67 61.11 70.12
Inp+Ans 75.33 53.89 67.98
Inp 66.00 50.90 70.81
R1 77.33 59.28 70.10
PT 29.33 38.32 66.16
Table 5: Results of generalization capacities of pruned models. Domain denotes the domain of pruning data.
Domain AIME24 LiveCodebench GPQA
Math 80.67 46.11 46.91
Coding 38.00 61.11 39.90
Science 64.64 53.59 70.12
Figure 6: Left: Comparison of performances with different number of shots for pruning. Right: Total Throughput across different numbers of experts.
5.3.4 Effect of Number of Pruning Demonstrations
In Section 3.2, we demonstrated that domain-relevant experts can be effectively identified using only a few demonstrations. Here, we further investigate the impact of the number of demonstrations on final performance. To do so, we sample varying numbers of demonstrations from the same distribution and prune half of the experts in each layer. The performance variations on AIME24 and LiveCodeBench are presented in Figure 6 (Left). When we utilize only a single sample for pruning, the selected experts are often influenced by the characteristics of that individual sample, thus resulting in lower performance. As we further increase the number of demonstrations, the performance rapidly rises, achieving comparable performances with the full model.
5.4 Analysis of Throughout
To evaluate the throughput of pruned models with different numbers of experts, we use the SGLang [17] package and measure performance under a maximum request concurrency of 32. We consider four settings: (1) 1K+1K: 1K input length and 1K output length; (2) 1K+4K: 1K input length and 4K output length; (3) 4K+1K: 4K input length and 1K output length; and (4) 4K+4K: 4K input length and 4K output length. For configurations with no more than 192 experts, we use a single 8$\times$×H800 GPU and double the measured throughput. For configurations with more than 192 experts, two 8$\times$×H800 GPUs are used. Figure 1 shows the scaling throughput for the 1K+1K setting, while Figure 6 (Right) presents results for the other settings.
We observe that reducing the number of experts significantly improves throughput, particularly when the model can be deployed on a single node. Compared to the full DeepSeek-R1 model, configurations with 128 and 64 experts achieve 2.99$\times$ and 4.33$\times$ higher throughput, respectively, under the 1K+1K setting. Under the 4K+4K setting, they reach 2.51$\times$ and 2.97$\times$ improvements. Furthermore, in long output scenarios with an equivalent total sequence length, the 128-expert configuration achieves a significantly higher acceleration ratio (2.91$\times$ under the 1K+4K setting, compared to 2.52$\times$ under the 4K+1K setting).
6 Related Work
6.1 Sparse MoE Architectures
Sparse MoE architectures aim to improve computational efficiency by activating only a small subset of experts for each input [18, 19, 4]. These methods focus on designing inherently efficient MoE architectures, leveraging conditional computation to scale up model capacity without linearly increasing inference cost. Driven by scaling laws, the parameter size of dense models has grown dramatically, leading to the emergence of even larger MoE models [20, 2, 21]. To further enhance the efficiency of MoE, DeepSpeed-MoE [22] proposes a pyramid-like architecture where the lower layers contain more experts. Similarly, DeepSeek-MoE [10] adopts a fine-grained expert design that reduces the parameter size per expert while increasing the total number of experts. In addition, another line of research focuses on transforming dense models into MoE architectures to reduce computational costs [23, 24]. By introducing sparsity into existing dense architectures, these methods aim to retain the performance of large models while benefiting from the conditional computation property of MoE. These studies primarily concentrate on architectural design. In contrast, another line of research aims to compress already-trained MoE models using post-hoc methods such as pruning and quantization, which we discuss in the next section.
6.2 MoE Compression Methods
While sparse MoE architectures improve computational efficiency by activating only a small subset of experts, they still incur significant memory overhead due to the need to store all expert parameters [25], posing a different set of challenges from existing dense LLM compression [26, 27]. The prior study [28] has observed that many experts contain redundant information, motivating efforts to further reduce memory usage by eliminating unnecessary experts. A common approach is to apply sparsification techniques to retain only the most important experts, based on carefully curated importance metrics. These metrics include the frequency [5] with which experts are selected during routing (i.e., “heavy hitter counts”), as well as analyses of expert gating scores and activation patterns [29, 5]. In addition to selection-based strategies, parameter sharing among experts can also be employed to alleviate the overall memory overhead [28, 30]. Some studies further assess expert importance by evaluating their impact on the overall model output, aiming to search for the optimal combination of experts [6, 7]. However, such methods involve exploring a large combinatorial search space, which leads to significant computational overhead and makes them impractical for large-scale MoE models. Additionally, another line of research leverages expert representations and router logits to merge similar experts into a single expert, thereby reducing the total number of experts [31, 32]. However, these methods introduce additional parameters and incur significant memory costs, particularly when performing multiple task-specific pruning. Different from them, our approach enables efficient pruning with just a single forward pass and eliminates the need for extra memory to store new models.
7 Conclusion
In this work, we investigated the domain specialization of experts in DeepSeek-R1. Our observations indicate that domain-specific experts play a crucial role in their respective domains and can be effectively identified with a few demonstrations. Building on these insights, we proposed a pruning strategy that leverages demonstrations from tasks within the same domain. Specifically, we introduced EASY-EP, a simple yet effective pruning method that combines output-aware expert importance assessment with expert-level token contribution estimation. Experimental results showed that our approach maintained comparable performance while utilizing only half of the experts in domain-specific settings and retains approximately $95\%$ of the original performance in mixed-domain pruning. We believe that our method can facilitate the deployment of large MoE models, particularly for efficiently handling a high volume of samples within the same domain.
Appendix A Experiment Details
As shown in the previous analysis in Section 3.2, datasets within the same domain can identify the important experts on other datasets. Thus, for different domains and tasks, we select one dataset as a calibration set, which are shown in Table 6. For mixed-domain pruning, we select the scores calculated on each 25-shot calibration set and average the normalized scores as the final scores of experts. Additionally, all the experiments are conducted in one 8×8$\times$8 × H200 GPU.
Table 6: Calibration Set of Each Domain
Domain Calibration Set
Math AIME 2023
Coding LiveCodeBench-V3
Science GPQA-Main
Agent Dev Set of AgentBench-OS
Appendix B Analysis of Perturbation-based Pruning Method
Previous studies  [6, 7] have proposed methods that utilize representation perturbation after expert pruning to determine which experts should be removed. In NAEE [6], identifying the optimal subset of experts requires $\operatorname{C}_{N}^{N^{\prime}}$ evaluations, where $N$ and $N^{\prime}$ represent the original and target numbers of experts, respectively. In CD-MoE [7], a greedy search algorithm selects the expert to retain based on minimal representation perturbation through a rolling mechanism, requiring $N(N+N^{\prime})/2$ evaluations. For DeepSeek-R1, which has 256 experts and a target expert count of 128, these methods require over $10^{75}$ and $24768$ evaluations per layer, respectively. Thus, the perturbation-based methods are not affordable for MoE models with a large number of experts. Conversely, our method only requires one forward operation to identify critical experts, which is cheaper and more suitable for large MoE models.